OpenStudy (anonymous):

Ca someone tell me how to do this 20. Solve 3/x + 5/x= -4 A. x=-1/2 B. x=2 C. x=1/2 D.x=-2

OpenStudy (anonymous):

you have to isolate x, or, get x by itself. currently x is in the denominator of two terms. you must: 1. group the denominator (Common Denominators) 2. Bring the x to the numerator (cross multiplication or reciprocals) 3. clear any coefficients

OpenStudy (anonymous):

Don't get caught up with x being in the denominator. This is just like elementary math. Fractions with the same denominator can be added

OpenStudy (campbell_st):

well you have \[\frac{3}{x} + \frac{5}{x} = -\] your fractions have the same denominator so add the numerators \[\frac{8}{x} = -4\] multiply both sides by x \[8 = -4x\] just solve for x

OpenStudy (anonymous):

i got x=2

OpenStudy (campbell_st):

ummm I think there is a problem... 8/(-4) is not 2

OpenStudy (anonymous):

its -2

OpenStudy (campbell_st):

thats correct

OpenStudy (anonymous):

Thank you